OpenStudy (anonymous):

Word Problems i need help i'm really stuck. :( Two ships start toward each other at the same time from 2 islands that are 264 km apart. One ship travels 4 km/h faster than the other ship. If they meet in 6 hours, how fast is each ship traveling?

OpenStudy (anonymous):

distance = speed x time going toward each other closes the distance additively 264 km = (v1 + v2) (6hrs) v2 = v1 + 4 264 km = (2 v1 + 4) (6 h) solve for v1 and thus also for v2.